### v1

# Potential Energy Curves of Hydrogenic Halides HX(Cl,Br) and i.DMFT Method

Abstract
Comparison of ab initio calculations in i.DMFT Method by Di Liu et al. (2025) with benchmark potential curves for HX(Cl,Br) halides shows their inaccuracy in domain around equilibrium and wrong behavior in Van der Waals region of large distances.
Recently, the ground state potential curves for diatomic molecules HF/DF/TF [1], HCl/DCl/TCl [2], HBr/DBr/TBr and HI/DI/TI [3] were constructed analytically with relative accuracy $\sim 10^{-4}$ (or four significant digits (s.d.)). This was based on the interpolation of perturbation theory at small internuclear distances $R$ with multipole expansion at large internuclear distances (Van der Waals regime) by using the a two-point Padé approximation
$E_{d}(R)\ =\ \frac{1}{R}\ \mbox{Pade}[N/N+5](R)\ \equiv\ \frac{1}{R}\ \frac{P_{N}}{Q_{N+5}}\ ,$ (1)
which is ratio of two polynomials, where $N=5$ was chosen. Parameters of the employed Padé approximant $\mbox{Pade}[5/10]$ were fixed by requirement to reproduce exactly the first three theoretically-known coefficients of perturbation theory at $R\rightarrow 0$ and two theoretically-known coefficients at $R\rightarrow\infty$ expansions as well as three spectroscopic, experimentally-known parameters/constants. The eight remaining free parameters of $E_{d}(R)$ (1) were fixed by describing accurately a four low-lying vibrational energies of HF, HCl, HBr and HI molecules with relative accuracy $\sim 10^{-4}$ (or four s.d.), respectively, via nuclear (two-body) Schrödinger equation by adjusting the corresponding eight turning points. Quality of obtained potential curves was checked by solving nuclear (two-body) radial Schrödinger equation (without breakdown functions!)  [^1] Hence, without modification of kinetic energy, potential energy and centrifugal potential with $E_{d}(R)$, see (1) at $N=5$, taken as a potential. It led to a description of all available experimental rovibrational energies with relative accuracy $\sim 10^{-4}$ (or four s.d.) for all above mentioned molecules and its isotopomers collected in [4]. This result was in agreement with domain of applicability of the Born-Oppenheimer approximation [5], where mass corrections $\sim m_{e}/M_{nuclei}$, QED corrections $\sim\alpha^{2}$, relativistic corrections $\sim v^{2}/c^{2}$ - are all of the order $\lesssim 10^{-4}$ - can be neglected. Hence, the obtained potential curves are benchmark curves written in the form (1) inside of the Born-Oppenheimer approximation. Obtained potential curves were in agreement with turning points positions, found in [4], on the level of 3-4 s.d.
Solving the electronic Schrödinger equation in order to find the potential curves (equivalently, their eigenvalues) is usually the subject of the so-called ab initio calculations, which is difficult problem per se, even for two center case, even for a few electron case, especially in evaluation of the rate of convergence of employed method and the accuracy of obtained results. In fact, the only real way of evaluation at present is by making a comparison with firmly established results earlier. In this Article we compare the results by Di Liu et al. [6], see Figs. 3-4 therein, based on i.DMFT method  [^2] We thank Dr. Di Liu for mailing us the numerical results on the potential curves for HCl and HBr diatomic molecules., with ground state potential curves for HCl and HBr molecules obtained in [2-3], see Tables 1-2 below, based on the potential (1) at $N=5$.
Table 1: Electronic ground state $X^{1}\Sigma^{+}$ of the diatomic molecule HCl: potential curve (in Hartree) vs. internuclear distance $R$ with the equilibrium at $R_{eq}=2.408542$ a.u. Second column from Di Liu et al.[6], third one from fit (1) [2].
$R$ (a.u.) Di Liu [6] Fit (1) [2]
0.944863 2.228391 2.050960
1.133836 1.170356 1.093860
1.511781 0.214895 0.201806
2.078699 -0.143275 -0.144284
2.408645 -0.169691 -0.169600
3.023562 -0.132582 -0.133642
4.062911 -0.049204 -0.051095
5.007774 -0.013979 -0.013794
6.047124 -0.004411 -0.002485
7.086473 -0.002508 -0.000504
7.464418 -0.002233 -0.000304
8.031336 -0.001956 -0.000154
8.598254 -0.001764 -0.000085
9.070685 -0.001639 -0.000055
9.448631 -0.001554 -0.000040
10.015548 -0.001446 -0.000026
10.582466 -0.001353 -0.000018
11.054898 -0.001284 -0.000013
12.094247 -0.001150 -7.392080E(-6)
12.566679 -0.001096 -5.807060E(-6)
13.039110 -0.001048 -4.616310E(-6)
13.511542 -0.001004 -3.707980E(-6)
14.078460 -0.000957 -2.885170E(-6)
14.550891 -0.000920 -2.361760E(-6)
15.023323 -0.000886 -1.947420E(-6)
15.495754 -0.000853 -1.616460E(-6)
16.062672 -0.000814 -1.303060E(-6)
16.535104 -0.000784 -1.095530E(-6)
17.007535 -0.000754 -9.258130E(-7)
17.952398 -0.000700 -6.705880E(-7)
18.897261 -0.000650 -4.940380E(-7)
37.794523 -0.000225 -7.934300E(-9)
56.691784 -0.000097 -7.012550E(-10)
75.589045 -0.000036 -1.251100E(-10)
[…]
Table 2: Electronic ground state $X^{1}\Sigma^{+}$ of the diatomic molecule HBr: potential curve (in Hartree) vs. internuclear distance $R$ with the equilibrium at $R_{eq}=2.67288$ a.u. Second column from Di Liu et al.[6], third one from fit (1) [3].
$R$ (a.u.) D.Liu [6] Fit (1) [3]
0.944863 3.157199 2.844940
1.039349 2.393633 2.183360
1.606267 0.349469 0.327671
2.078699 -0.054857 -0.059769
2.456644 -0.135701 -0.136453
2.645617 -0.143933 -0.143925
3.023562 -0.132205 -0.132035
3.590480 -0.090917 -0.091634
4.062911 -0.057631 -0.058866
4.535343 -0.032510 -0.033843
5.007774 -0.016445 -0.017690
5.669178 -0.005742 -0.006566
6.047124 -0.003210 -0.003725
6.614041 -0.001516 -0.001659
7.086473 -0.000957 -0.000891
7.653391 -0.000681 -0.000452
8.031336 -0.000599 -0.000299
8.598254 -0.000537 -0.000170
9.070685 -0.000507 -0.000110
9.448631 -0.000488 -0.000080
10.015548 -0.000465 -0.000052
10.582466 -0.000450 -0.000034
11.054898 -0.000444 -0.000025
11.527329 -0.000440 -0.000019
12.094247 -0.000436 -0.000013
12.566679 -0.000431 -0.000010
13.039110 -0.000424 -8.056090E(-6)
14.078460 -0.000407 -4.852980E(-6)
15.023323 -0.000393 -3.183680E(-6)
17.007535 -0.000353 -1.444820E(-6)
18.897261 -0.000308 -7.470350E(-7)
37.794523 -0.000105 -1.090780E(-8)
56.691784 -0.000046 -9.492540E(-10)
75.589045 -0.000017 -1.684900E(-10)
[…]
One can see that systematically the i.DMFT Method does not reproduce four s.d. in benchmark potential curves (1), thus, deviates from turning points obtained in [4], except for a domain in vicinity of minimum, where there is an agreement in 3 s.d. Absolute deviation becomes truly significant beyond equilibrium internuclear distances, at small and especially at large distances. At R>7R>7a.u. for HCl molecule and at R>9R>9a.u. for HBr molecule the deviation is growing with the increase of $R$ and can reach several orders of magnitude for a large $R$, see Tables 1-2, in particular, at $R\sim 75$ a.u. it reaches five orders of magnitude! Hence, it indicates that the behavior of potential curves obtained in [6] in Van der Waals regime does not agree with the multipole expansion. It implies that by taking points found in the i.DMFT Method as basic ones on some potential curve will lead to inaccurate rovibrational spectra, even for low-lying states. It is confirmed by making a comparison of the vibrational energies $E_{\nu 0}$ obtained in [6] (see Supporting information therein) with results from [2-4]: one can see that for small $\nu$ three s.d. are reproduced only, then for larger $\nu<17$ - one-two s.d. and for $\nu=17-20$ even single s.d. is not reproduced! It indicates to shortcomings of the i.DMFT Method.
Acknowledgments
This work is partially supported by DGAPA grant IN113025 (Mexico). We are grateful to Prof. Andrew Orr-Ewing, the JPCA Editor, for suggestions which led to the improvement and the clarity of the presentation.