### v2

# Potential Energy Curves of Hydrogenic Halides HX(Cl,Br) and i-DMFT Method

Abstract
A comparison of the ab initio calculations using the i-DMFT Method by Di Liu et al. (2025) with benchmark potential curves for HX(Cl,Br) halides shows their inaccuracy in the domain around equilibrium - they do not reproduce quantitatively the results of the Born-Oppenheimer approximation - and also they predict a qualitatively wrong behavior in the Van der Waals region of large distances, thus, contradict the multipole expansion.
Recently, the ground state potential curves for diatomic molecules HF/DF/TF [1], HCl/DCl/TCl [2], HBr/DBr/TBr and HI/DI/TI [3] were constructed analytically with a relative accuracy $\sim 10^{-4}$ (or four significant digits (s.d.)). This was based on the interpolation of perturbation theory at small internuclear distances $R$ with the multipole expansion at large internuclear distances (Van der Waals regime) by using a two-point Padé approximation
$E_{d}(R)\ =\ \frac{1}{R}\ \mbox{Pade}[N/N+5](R)\ \equiv\ \frac{1}{R}\ \frac{P_{N}}{Q_{N+5}}\ ,$ (1)
which is ratio of two polynomials, where $N=5$ was chosen. The parameters of the employed Padé approximant $\mbox{Pade}[5/10]$ were fixed by requiring the exact reproduction of the first three theoretically-known coefficients of perturbation theory at $R\rightarrow 0$ and the two theoretically-known coefficients at $R\rightarrow\infty$ expansions as well as three spectroscopic, experimentally-known parameters/constants. The eight remaining free parameters of $E_{d}(R)$ (1) were fixed by describing accurately the four low-lying vibrational energies of the HF, HCl, HBr and HI molecules with relative accuracy $\sim 10^{-4}$ (or four s.d.), respectively, via the nuclear (two-body) Schrödinger equation by adjusting the corresponding eight turning points. The quality of the obtained potential curves was checked by solving the nuclear (two-body) radial Schrödinger equation (without breakdown functions!)  [^1] Hence, without modification of the kinetic energy, potential energy and centrifugal potential. with $E_{d}(R)$, see (1) at $N=5$, taken as a potential. This led to a description of all available experimental rovibrational energies with relative accuracy $\sim 10^{-4}$ (or four s.d.) for all above mentioned molecules and its isotopomers collected in [4]. This result was in agreement with the domain of validity of the Born-Oppenheimer approximation [5], where mass corrections $\sim m_{e}/M_{nuclei}$, QED corrections $\sim\alpha^{2}$, relativistic corrections $\sim v^{2}/c^{2}$ - all of the order $\lesssim 10^{-4}$ - can be neglected. Hence, the obtained potential curves are benchmark curves written in the form (1) inside of the Born-Oppenheimer approximation. It must be emphasized that the obtained potential curves were in agreement with the turning point positions, found in [4], on the level of 3-4 s.d. Hence, the potential energy curves (1) at $N=5$ approximate the lowest eigenvalue of the electronic Hamiltonian with relative accuracy $\sim 10^{-4}$.
Solving the electronic Schrödinger equation in order to find the potential energy curves (equivalently, their eigenvalues) is usually the subject of the so-called ab initio calculations, which is a difficult problem per se, even for the two center case, even for a few electron case, especially in the evaluation of the rate of convergence of the employed method and the accuracy of obtained results. In fact, the only real way of evaluation (at present) is by making a comparison with the firmly established results obtained earlier. In this short Article we compare the results by Di Liu et al. [6], see Figs. 3-4 therein, based on the recently-proposed i-DMFT method  [^2] We thank Dr. Di Liu for mailing us the numerical results on the potential curves for HCl and HBr diatomic molecules., with the ground state potential curves for HCl and HBr molecules obtained in [2-3], see Tables 1-2 below, based on the potential (1) at $N=5$.
Table 1: Electronic ground state $X^{1}\Sigma^{+}$ of the diatomic molecule HCl: potential curve (in Hartree) vs. internuclear distance $R$ with the equilibrium at $R_{eq}=2.408542$ a.u. Second column from Di Liu et al.[6], third one from fit (1) [2].
$R$ (a.u.) Di Liu [6] Fit (1) [2]
0.944863 2.228391 2.050960
1.133836 1.170356 1.093860
1.511781 0.214895 0.201806
2.078699 -0.143275 -0.144284
2.408645 -0.169691 -0.169600
3.023562 -0.132582 -0.133642
4.062911 -0.049204 -0.051095
5.007774 -0.013979 -0.013794
6.047124 -0.004411 -0.002485
7.086473 -0.002508 -0.000504
7.464418 -0.002233 -0.000304
8.031336 -0.001956 -0.000154
8.598254 -0.001764 -0.000085
9.070685 -0.001639 -0.000055
9.448631 -0.001554 -0.000040
10.015548 -0.001446 -0.000026
10.582466 -0.001353 -0.000018
11.054898 -0.001284 -0.000013
12.094247 -0.001150 -7.392080E(-6)
12.566679 -0.001096 -5.807060E(-6)
13.039110 -0.001048 -4.616310E(-6)
13.511542 -0.001004 -3.707980E(-6)
14.078460 -0.000957 -2.885170E(-6)
14.550891 -0.000920 -2.361760E(-6)
15.023323 -0.000886 -1.947420E(-6)
15.495754 -0.000853 -1.616460E(-6)
16.062672 -0.000814 -1.303060E(-6)
16.535104 -0.000784 -1.095530E(-6)
17.007535 -0.000754 -9.258130E(-7)
17.952398 -0.000700 -6.705880E(-7)
18.897261 -0.000650 -4.940380E(-7)
37.794523 -0.000225 -7.934300E(-9)
56.691784 -0.000097 -7.012550E(-10)
75.589045 -0.000036 -1.251100E(-10)
[…]
Table 2: Electronic ground state $X^{1}\Sigma^{+}$ of the diatomic molecule HBr: potential curve (in Hartree) vs. internuclear distance $R$ with the equilibrium at $R_{eq}=2.67288$ a.u. Second column from Di Liu et al.[6], third one from fit (1) [3].
$R$ (a.u.) D.Liu [6] Fit (1) [3]
0.944863 3.157199 2.844940
1.039349 2.393633 2.183360
1.606267 0.349469 0.327671
2.078699 -0.054857 -0.059769
2.456644 -0.135701 -0.136453
2.645617 -0.143933 -0.143925
3.023562 -0.132205 -0.132035
3.590480 -0.090917 -0.091634
4.062911 -0.057631 -0.058866
4.535343 -0.032510 -0.033843
5.007774 -0.016445 -0.017690
5.669178 -0.005742 -0.006566
6.047124 -0.003210 -0.003725
6.614041 -0.001516 -0.001659
7.086473 -0.000957 -0.000891
7.653391 -0.000681 -0.000452
8.031336 -0.000599 -0.000299
8.598254 -0.000537 -0.000170
9.070685 -0.000507 -0.000110
9.448631 -0.000488 -0.000080
10.015548 -0.000465 -0.000052
10.582466 -0.000450 -0.000034
11.054898 -0.000444 -0.000025
11.527329 -0.000440 -0.000019
12.094247 -0.000436 -0.000013
12.566679 -0.000431 -0.000010
13.039110 -0.000424 -8.056090E(-6)
14.078460 -0.000407 -4.852980E(-6)
15.023323 -0.000393 -3.183680E(-6)
17.007535 -0.000353 -1.444820E(-6)
18.897261 -0.000308 -7.470350E(-7)
37.794523 -0.000105 -1.090780E(-8)
56.691784 -0.000046 -9.492540E(-10)
75.589045 -0.000017 -1.684900E(-10)
[…]
One can see that systematically the i-DMFT Method does not reproduce four s.d. in the benchmark potential energy curves (1), thus, it deviates from the turning points obtained in [4], except for a domain in the vicinity of the minimum, where there is an agreement in 3 s.d. The absolute deviation becomes truly significant beyond equilibrium internuclear distances, at small and especially at large distances. In particular, at R>7R>7a.u. for the HCl molecule and at R>9R>9a.u. for the HBr molecule the deviation is growing with the increase of $R$ and can reach several orders of magnitude for a large $R$, see Tables 1-2; in particular, at $R\sim 50-75$ a.u. it reaches five orders of magnitude! Hence, this indicates that the behavior of the potential energy curves obtained in [6] in the Van der Waals regime does not agree with the multipole expansion. This implies that by taking the turning points found using the i-DMFT Method as basic ones and placing them on some potential curve will lead to inaccurate rovibrational spectra, even for low-lying states. It is confirmed by making a comparison of the vibrational energies $E_{\nu 0}$ obtained in [6] (see Supporting information therein) with the results from [2-4] 333which are in agreement with experimental vibrational spectra in, at least, four s.d.: one can see that for small $\nu$ three s.d. are reproduced only (as expected), then for larger $\nu<17$ - one-two s.d. while for $\nu=17-20$ not even single s.d. is reproduced! This indicates shortcomings of the i-DMFT Method. The question, of how intrinsic to the method used these shortcomings are, needs to be investigated and clarified before drawing conclusions about the heuristic value of the i-DMFT Method.
Acknowledgments
This work is partially supported by DGAPA grant IN113025 (Mexico). We are grateful to J. Karwowski and A. Orr-Ewing, the JPCA Editor, for suggestions which led to the improvement and the further clarity of the presentation.